OpenStudy (anonymous):

under the QUADRATIC EQUATION(Roots) If a triangle has an area of 56 sq. inches, with base of 9 inches longer that its altitude, find the length of its base.

OpenStudy (anonymous):

if altitude means height then this is what i'd do set the altitude as x and base as x+9 then the area is 1/2 base*hight, (1/2)*(x*(x+9)=56) x^2+9x=108 x^2+9x-108=0 now solve for the quadratic and you get \[x=\frac{9\pm \sqrt{513}}{2}\]